# FlickerFusion: Intra-trajectory Domain Generalizing Multi-Agent RL

### v1

@@ 5 Empirical Study
ODIS (Zhang et al., 2023) -257.1±${}_{\pm\text{38.2}}$376.5±${}_{\pm\text{84.6}}$3532.8±${}_{\pm\text{2994.9}}$2982.1±${}_{\pm\text{2074.4}}$1992.0±${}_{\pm\text{2331.4}}$1723.4±${}_{\pm\text{1817.3}}$
@@
ODIS (Zhang et al., 2023) -1449.3 ±${}_{\pm\text{80.1}}$1442.8 ±${}_{\pm\text{278.6}}$56.8 ±${}_{\pm\text{37.7}}$188.7 ±${}_{\pm\text{105.2}}$667.5 ±${}_{\pm\text{99.5}}$657.8 ±${}_{\pm\text{209.7}}$
@@
We examine the empirical performance of FlickerFusion against 11 relevant baselines. The baselines are divided into 3 categories: (i) MARL backbone methods, (ii) MARL backbone methods with model-agnostic domain generalization methods, and (iii) recent relevant MARL methods. (i) The two backbone archiecture we use are QMIX-MLP (Rashid et al., 2018) and QMIX-Attention (Iqbal et al., 2021). (ii) To ensure that our method is most competitive vis-à-vis domain generalization, we implement four model-agnostic domain generalization methods on top of the best performing backbone, QMIX-Attention. These are MLDG (Li et al., 2018), SMLDG (Li et al., 2020), DG-MAML (Wang et al., 2021a), and Meta DotProd (Kedia et al., 2021). More recent methods are highly domain- and model-specific, leading to logically flawed implementations if applied to our problem setting. We are the first to enact such rigor in empirically examining model-agnostic domain generalization methods in the MARL problem setting. (iii) The remaining five baselines are most relevant to our problem setting. UPDeT (Hu et al., 2021), REFIL (Iqbal et al., 2021), CAMA (Shao et al., 2023), and ODIS (Zhang et al., 2023) are representative methods that focus on dynamic team composition. CAMA and ODIS specifically study zero-shot OOD generalization and show state-of-the-art results. Finally, we include the most recently proposed state-of-the-art general MARL method, ACORM (Hu et al., 2024).
While related, we do not include Wang et al. (2023a) and Tian et al. (2023). Wang et al. (2023a)’s implementation is complex, closed-source, and only empirically examined with a PPO (Schulman et al., 2017) backbone. On-policy RL backbones are notoriously expensive to train for MARL, resulting in poor scalability (Papoudakis et al., 2021). Similarly, Tian et al. (2023)’s implementation is complex, closed-source, with only two baselines and no study against their backbone architecture.
@@
Figure 5: Additional empirical study visualizations. Top-left is the box-and-whisker plot uncertainty distributions across methods. The remainder are reward curves of Repel (OOD1), demonstrating that FlickerFusion (red, dark blue) not only performs well at 3M steps, but also converges faster.
@@ 6 Discussion and Takeaways
(1) Novel universal inductive bias injection.
@@
(2) Potential in non-attention backbones.
Unlike other MARL generalization methods that only work with an attention backbone (Shao et al., 2023; Zhang et al., 2023), FlickerFusion works exceptionally well under MLP and attention architectures. Surprisingly, FlickerFusion-MLP often beats all other existing attention-based methods, even with smaller model (parameter) size (Table 1). Further, during hyperparameter tuning, all else equal, we find no meaningful relationship between model size (parameters) and performance. This echoes other domain literature, which shows that attention may not always be the most optimal design choice (Liu et al., 2021; Zeng et al., 2023).
(3) No reward-uncertainty trade-off.
@@
(4) Online over offline data.
@@
(5) Naïve model-agnostic approaches are ill-advised for MARL.
@@ Acknowledgments
Let us now convert the above in-expectation guarantee to a high-probability guarantee via McDiarmid’s inequality (McDiarmid, 1989). Define $s_{i,a}=\mathds{1}[i\in D_{a}]\in\{0,1\}$ for $(i,a)\in[N_{\ell}]\times[N_{A}]$, which is independent across $a$. Consider the function $f$ defined as $f(\{s_{i,a}\}_{a\in[N_{A}]}):=\left|\frac{1}{N_{A}}\sum_{a\in[N_{A}]}s_{i,a}-\frac{\Delta_{\ell}}{N_{\ell}}\right|$. This satisfies the bounded difference property with constant $\frac{1}{N_{A}}$. Thus, we have that for any $\varepsilon>0$,
${\mathbb{P}}\left(\sum_{i\in[N_{\ell}]}f(\{s_{i,a}\}_{a\in[N_{A}]})-\mathbb{E}[f(\{s_{i,a}\}_{a\in[N_{A}]})]\geq\varepsilon\right)$
$\leq{\mathbb{P}}\left(\exists i\in[N_{\ell}]\ :\ f(\{s_{i,a}\}_{a\in[N_{A}]})-\mathbb{E}[f(\{s_{i,a}\}_{a\in[N_{A}]})]\geq\frac{\varepsilon}{N_{\ell}}\right)$
$\leq\sum_{i\in[N_{\ell}]}{\mathbb{P}}\left(f(\{s_{i,a}\}_{a\in[N_{A}]})-\mathbb{E}[f(\{s_{i,a}\}_{a\in[N_{A}]})]\geq\frac{\varepsilon}{N_{\ell}}\right)$ (union bound)
$\leq N_{\ell}\exp\left(-\frac{2(\varepsilon/N_{\ell})^{2}}{(1/N_{A})^{2}}\right)$ (McDiarmid’s inequality)
$=N_{\ell}\exp\left(-2\varepsilon^{2}\left(\frac{N_{A}}{N_{\ell}}\right)^{2}\right).$
All in all, we have that
$1-N_{\ell}\exp\left(-2\varepsilon^{2}\left(\frac{N_{A}}{N_{\ell}}\right)^{2}\right)$ $\leq{\mathbb{P}}\left(\sum_{i\in[N_{\ell}]}\left|\frac{d(i)}{N_{A}}-\frac{\Delta_{\ell}}{N_{\ell}}\right|\leq\varepsilon+\mathbb{E}\left[\sum_{i\in[N_{\ell}]}\left|\frac{d(i)}{N_{A}}-\frac{\Delta_{\ell}}{N_{\ell}}\right|\right]\right)$
$\leq{\mathbb{P}}\left(\sum_{i\in[N_{\ell}]}\left|\frac{d(i)}{N_{A}}-\frac{\Delta_{\ell}}{N_{\ell}}\right|\leq\varepsilon+\sqrt{\frac{\Delta_{\ell}(N_{\ell}-\Delta_{\ell})}{N_{A}}}\right).$
Reparametrizing $\delta=N_{\ell}\exp\left(-2\varepsilon^{2}\left(\frac{N_{A}}{N_{\ell}}\right)^{2}\right)$ gives the desired result. ∎

### v4

@@ 5 Empirical Study
ODIS (Zhang et al., 2023a) -257.1±${}_{\pm\text{38.2}}$376.5±${}_{\pm\text{84.6}}$3532.8±${}_{\pm\text{2994.9}}$2982.1±${}_{\pm\text{2074.4}}$1992.0±${}_{\pm\text{2331.4}}$1723.4±${}_{\pm\text{1817.3}}$
@@
ODIS (Zhang et al., 2023a) -1449.3 ±${}_{\pm\text{80.1}}$1442.8 ±${}_{\pm\text{278.6}}$56.8 ±${}_{\pm\text{37.7}}$188.7 ±${}_{\pm\text{105.2}}$667.5 ±${}_{\pm\text{99.5}}$657.8 ±${}_{\pm\text{209.7}}$
@@
We examine the empirical performance of FlickerFusion against 11 relevant baselines. The baselines are divided into 3 categories: (i) MARL backbone methods, (ii) MARL backbone methods with model-agnostic domain generalization methods, and (iii) recent relevant MARL methods. (i) The two backbone architecture we use are QMIX-MLP (Rashid et al., 2018) and QMIX-Attention (Iqbal et al., 2021). (ii) To ensure that our method is most competitive vis-à-vis domain generalization, we implement four model-agnostic domain generalization methods on top of the best performing backbone, QMIX-Attention. These are MLDG (Li et al., 2018), SMLDG (Li et al., 2020), DG-MAML (Wang et al., 2021a), and Meta DotProd (Kedia et al., 2021). More recent methods are highly domain- and model-specific, leading to logically flawed implementations if applied to our problem setting. We are the first to enact such rigor in empirically examining model-agnostic domain generalization methods in the MARL problem setting. (iii) The remaining five baselines are most relevant to our problem setting. UPDeT (Hu et al., 2021), REFIL (Iqbal et al., 2021), CAMA (Shao et al., 2023), and ODIS (Zhang et al., 2023a) are representative methods that focus on dynamic team composition. CAMA and ODIS specifically study zero-shot OOD generalization and show state-of-the-art results. Finally, we include the most recently proposed state-of-the-art general MARL method, ACORM (Hu et al., 2024).
While related, we do not include Wang et al. (2023a) and Tian et al. (2023). Wang et al. (2023a)’s implementation is complex, closed-source, and only empirically examined with a PPO (Schulman et al., 2017) backbone. On-policy RL backbones are notoriously expensive to train for MARL, resulting in poor scalability (Papoudakis et al., 2021). Similarly, Tian et al. (2023)’s implementation is complex, closed-source, with only two baselines and no study against their backbone architecture. For readers that are interested in improved generalization through multi-task learning, they can refer to Qin et al. (2024).
@@
Figure 5: Additional empirical study visualizations. Top-left is the box-and-whisker plot uncertainty distributions across methods. The remainder are reward curves of Repel (OOD1), demonstrating that FlickerFusion (dark blue, red) not only performs well at 3M steps, but also converges faster. The shaded regions indicate quartile 1 to 3—complementing the top-left standard deviation statistics.
@@ 6 Discussion and Takeaways
Novel universal inductive bias injection.
@@
Potential in non-attention backbones.
Unlike other MARL generalization methods that only work with an attention backbone (Shao et al., 2023; Zhang et al., 2023a), FlickerFusion works exceptionally well under MLP and attention architectures. Surprisingly, FlickerFusion-MLP often beats all other existing attention-based methods, even with smaller model (parameter) size (Table 1). Further, during hyperparameter tuning, all else equal, we find no meaningful relationship between model size (parameters) and performance. This echoes other domain literature, which shows that attention may not always be the most optimal design choice (Liu et al., 2021; Zeng et al., 2023).
No reward-uncertainty trade-off.
@@
Online over offline data.
@@
Naïve model-agnostic approaches are ill-advised for MARL.